# Stability of Poiseuille Flow of Navier-Stokes Equations on $\mathbb{R}^{2}$

### v1

@@ 3.6 Bound on $T_{5}$
$I_{3}\lesssim\left(\int\langle k\rangle^{-2m}(\gamma_{k}\nu)^{-1}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle k\rangle^{2m}\gamma_{k}\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\lesssim\nu^{-\frac{1}{6}}\mathcal{D}^{\frac{1}{2}}.$
@@
$|T_{5,HL,H^{\prime\prime}}|\lesssim\mathcal{E}^{\frac{1}{2}}\nu^{-\frac{1}{3}}\mathcal{D}^{\frac{1}{2}}\nu^{-\frac{1}{6}}\mathcal{D}^{\frac{1}{2}}=\nu^{-\frac{1}{2}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
@@
$|T_{5,HL,L^{\prime\prime}}|$ $\lesssim\iint 1_{|k^{\prime}|\lesssim\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|y\omega_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}|k-k^{\prime}|^{\frac{1}{2}}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}|k^{\prime}|^{-\frac{1}{2}}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}\|y\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|y\nabla_{k}\omega_{k}\|_{2}\mathrm{d}k.$
For the last term, using Hölder’s inequality,
$\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|y\nabla_{k}\omega_{k}\|_{2}\mathrm{d}k\leq\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\mathrm{d}k\sup_{k}\|y\nabla_{k}\omega_{k}\|_{2}\lesssim\nu^{-\frac{1}{6}}(\nu^{-1}\mathcal{D})^{\frac{1}{2}}=\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}.$
Combining Lemma 3.5, we have
$|T_{5,HL,L^{\prime\prime}}|\lesssim(\nu^{-\frac{2}{3}}\mathcal{E})^{\frac{1}{2}}(\nu^{-\frac{1}{3}}\mathcal{D})^{\frac{1}{2}}\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}}=\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
@@ 3.7 Bound on $T_{6}$
$T_{6}=-2c_{\gamma}\int\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\psi_{k},NL_{k}\rangle\mathrm{d}k.$
@@
and correspondingly we write $T_{6}\eqqcolon T_{6,LH,H}+T_{6,LH,L}+T_{6,HL,H}+T_{6,HL,L,LH}+T_{6,HL,L,HL}$. For $T_{6,LH,H}$, by Cauchy-Schwarz inequality in $y$, Young’s convolution inequality and Lemma 3.4,
$|T_{6,LH,H}|$ $\lesssim\iint 1_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\gamma_{k}\|\psi_{k}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k.$
@@
$\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k$ $=\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\beta_{k}|k|(\nu^{-\frac{1}{2}}|k|^{\frac{1}{2}})^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k$
$\lesssim\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-1}(\nu^{-\frac{1}{3}})^{-2}\beta_{k}|k|^{2}\|k\psi_{k}\|_{2}^{2}\mathrm{d}k$
$\lesssim\nu^{-\frac{1}{3}}\mathcal{D},$
@@
$|T_{6,LH,H}|\lesssim(\nu^{-\frac{1}{3}}\mathcal{D})^{\frac{1}{2}}(\nu^{-1}\mathcal{D})^{\frac{1}{2}}\mathcal{E}=\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
We treat $T_{6,LH,L}$ by separating it into two parts,
$T_{6,LH,L}$ $=2c_{\gamma}\int 1_{|k-k^{\prime}|\leq\frac{|k|}{2}}1_{|k|<\nu^{-\frac{1}{3}}}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\left\langle\psi_{k},\int i(k-k^{\prime})\psi_{k-k^{\prime}}\partial_{y}\omega_{k^{\prime}}\mathrm{d}k^{\prime}\right\rangle\mathrm{d}k$
$-2c_{\gamma}\int 1_{|k-k^{\prime}|\leq\frac{|k|}{2}}1_{|k|<\nu^{-\frac{1}{3}}}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\left\langle\psi_{k},\int\partial_{y}\psi_{k-k^{\prime}}ik^{\prime}\omega_{k^{\prime}}\mathrm{d}k^{\prime}\right\rangle\mathrm{d}k$
@@
$T_{6,LH,L}^{x}=-2c_{\gamma}\iint 1_{|k-k^{\prime}|\leq\frac{|k|}{2}}1_{|k|<\nu^{-\frac{1}{3}}}\frac{\langle c\lambda_{k}t\rangle^{2J}}{M_{k}(t)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\partial_{y}\psi_{k}\overline{\psi_{k-k^{\prime}}}+\psi_{k}\partial_{y}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}.$
By the triangle inequality and Cauchy-Schwarz inequality in $y$,
$|T_{6,LH,L}^{x}|$ $\lesssim\iint\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}(\|\partial_{y}\psi_{k}\|_{2}|k-k^{\prime}|\|\psi_{k-k^{\prime}}\|_{\infty}+|k|\|\psi_{k}\|_{2}\|\partial_{y}\psi_{k-k^{\prime}}\|_{\infty})\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime},$
where we have used $|k-k^{\prime}|\lesssim|k^{\prime}|$, since $|k-k^{\prime}|\leq\frac{|k|}{2}$. Then using Young’s convolution inequality and Lemma 3.4, we obtain
$|T_{6,LH,L}^{x}|\lesssim\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
@@
$|T_{6,LH,L}^{y}|\lesssim\iint\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\lesssim\nu^{-\frac{7}{6}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$
@@
$|T_{6,HL,H}|$ $\lesssim\iint 1_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{J}\langle k\rangle^{m}\gamma_{k}\|\psi_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}t\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{2}\|\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}$
$\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k.$
@@
$\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}t\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k\leq\nu^{-\frac{1}{3}}\mathcal{D},$
which implies
$|T_{6,HL,H}|\lesssim(\nu^{-\frac{1}{3}}\mathcal{D})^{\frac{1}{2}}(\nu^{\frac{2}{3}}\mathcal{E})^{\frac{1}{2}}\nu^{-\frac{5}{6}}\mathcal{D}^{\frac{1}{2}}=\nu^{-\frac{2}{3}}\mathcal{D}\mathcal{E}^{\frac{1}{2}}.$

### v4

@@ 3.4 Bound on $T_{5}$
$|T_{5,LH,L}|\lesssim\nu^{-\frac{1}{2}}\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}\left(\int_{0}^{t}\widetilde{\mathcal{D}}\mathrm{d}s\right)^{\frac{1}{2}}\left(\int_{0}^{t}\left(\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
With the LH terms bounded, we now turn to estimate the HL terms, starting with $T_{5,HL,H^{\prime\prime}}$. We have $\gamma_{k}\lesssim\gamma_{k-k^{\prime}}$, since $|k|\lesssim|k-k^{\prime}|$. Thus
$|T_{5,HL,H^{\prime\prime}}|$ $\lesssim\int_{0}^{t}\iint_{|k-k^{\prime}|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\gamma_{k}^{\frac{1}{2}}\|y\omega_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}\gamma_{k-k^{\prime}}^{\frac{1}{2}}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}\|y\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|y\nabla_{k}\omega_{k}\|_{2}\mathrm{d}k\mathrm{d}s$
$\eqqcolon\int_{0}^{t}I_{1}\cdot I_{2}\cdot I_{3}\mathrm{d}s.$
According to the definition of $\mathcal{E}$, we have $I_{1}\lesssim\mathcal{E}^{\frac{1}{2}}$. For $I_{2}$, we use Gagliardo-Nirenberg-Sobolev inequality, together with $\|k\nabla_{k}\psi_{k}\|_{2}\lesssim\|\Delta_{k}\psi_{k}\|_{2}=\|\omega_{k}\|_{2},\|\partial_{y}\nabla_{k}\psi_{k}\|\lesssim\|\Delta_{k}\psi_{k}\|_{2}=\|\omega_{k}\|_{2}$,
$I_{2}\lesssim\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}|k|^{-1}|k|\|\nabla_{k}\psi_{k}\|_{2}\|\partial_{y}\nabla_{k}\psi_{k}\|_{2}\mathrm{d}k\right)^{\frac{1}{2}}\lesssim\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}(\nu^{-\frac{1}{3}})^{-1}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\lesssim\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}}^{\frac{1}{2}}.$
Turning to $I_{3}$, by Hölder’s inequality and $\gamma_{k}\geq\nu^{-\frac{2}{3}}$, we obtain
$I_{3}\lesssim\left(\int_{\mathbb{R}}\langle k\rangle^{-2m}(\gamma_{k}\nu)^{-1}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle k\rangle^{2m}\gamma_{k}\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\lesssim\nu^{-\frac{1}{6}}\widetilde{\mathcal{D}}^{\frac{1}{2}}.$
@@
$|T_{5,HL,H^{\prime\prime}}|\lesssim\int_{0}^{t}\mathcal{E}^{\frac{1}{2}}\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}}^{\frac{1}{2}}\nu^{-\frac{1}{6}}\widetilde{\mathcal{D}}^{\frac{1}{2}}\mathrm{d}s\lesssim\nu^{-\frac{1}{2}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@
$|T_{5,HL,L^{\prime\prime}}|\!$ $\lesssim\!\int_{0}^{t}\!\iint_{|k^{\prime}|\lesssim\nu^{-\frac{1}{3}}}\!\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|y\omega_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}|k-k^{\prime}|^{\frac{1}{2}}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}|k^{\prime}|^{-\frac{1}{2}}\|y\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}\|y\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}|k|\|\nabla_{k}\psi_{k}\|_{\infty}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|y\nabla_{k}\omega_{k}\|_{2}\mathrm{d}k\mathrm{d}s$
$\lesssim\sup_{s\in[0,t]}(\nu^{-\frac{2}{3}}\mathcal{E}(s))^{\frac{1}{2}}\left(\int_{0}^{t}\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}}\mathrm{d}s\right)^{\frac{1}{2}}\left(\int_{0}^{t}\left(\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|y\nabla_{k}\omega_{k}\|_{2}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}},$
where we have used Lemma 3.6. For the last term, using Minkowski’s inequality for integrals,
$\left(\int_{0}^{t}\left(\int_{|k|\lesssim\nu^{-\frac{1}{3}}}|k|^{-\frac{1}{2}}\|y\nabla_{k}\omega_{k}\|_{2}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}$ $\lesssim\int_{|k|\lesssim\nu^{-\frac{1}{3}}}\left(\int_{0}^{t}\left(|k|^{-\frac{1}{2}}\|y\nabla_{k}\omega_{k}\|_{2}\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}\mathrm{d}k$
$\lesssim\int_{|k|\lesssim\nu^{-\frac{1}{3}}}\nu^{-\frac{1}{2}}|k|^{-\frac{1}{2}}\mathrm{d}k\cdot\sup_{k\in\mathbb{R}}\left(\int_{0}^{t}\nu\|y\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}s\right)^{\frac{1}{2}}$
$\lesssim\nu^{-\frac{2}{3}}\mathcal{D}^{\frac{1}{2}},$
which implies
$|T_{5,HL,L^{\prime\prime}}|\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@ 3.5 Bound on $T_{6}$
$T_{6}=-2c_{\gamma}\int_{0}^{t}\int_{\mathbb{R}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\psi_{k},NL_{k}\rangle\mathrm{d}k\mathrm{d}s.$
@@
and correspondingly we write $T_{6}\eqqcolon T_{6,LH,H}+T_{6,LH,L}+T_{6,HL,H}+T_{6,HL,L,LH}+T_{6,HL,L,HL}$. For $T_{6,LH,H}$, by Cauchy-Schwarz inequality in $y$ and Young’s convolution inequality,
$|T_{6,LH,H}|$ $\lesssim\int_{0}^{t}\iint_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\gamma_{k}\|\psi_{k}\|_{2}\langle c\lambda_{k^{\prime}}s\rangle^{J}\langle k^{\prime}\rangle^{m}\|\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s.$
@@
$\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k$ $=\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\beta_{k}|k|(\nu^{-\frac{1}{2}}|k|^{\frac{1}{2}})^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k$
$\lesssim\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-1}(\nu^{-\frac{1}{3}})^{-2}\beta_{k}|k|^{2}\|k\psi_{k}\|_{2}^{2}\mathrm{d}k$
$\lesssim\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}},$
@@
$|T_{6,LH,H}|\lesssim\int_{0}^{t}(\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}})^{\frac{1}{2}}(\nu^{-1}\widetilde{\mathcal{D}})^{\frac{1}{2}}\mathcal{E}^{\frac{1}{2}}\mathrm{d}s\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}},$
where we have used Lemma 3.4. We treat $T_{6,LH,L}$ by separating it into two parts,
$T_{6,LH,L}=$ $2c_{\gamma}\int_{0}^{t}\int_{|k|<\nu^{-\frac{1}{3}}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\left\langle\psi_{k},\int_{|k-k^{\prime}|\leq\frac{|k|}{2}}i(k-k^{\prime})\psi_{k-k^{\prime}}\partial_{y}\omega_{k^{\prime}}\mathrm{d}k^{\prime}\right\rangle\mathrm{d}k\mathrm{d}s$
$-2c_{\gamma}\int_{0}^{t}\int_{|k|<\nu^{-\frac{1}{3}}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\left\langle\psi_{k},\int_{|k-k^{\prime}|\leq\frac{|k|}{2}}\partial_{y}\psi_{k-k^{\prime}}ik^{\prime}\omega_{k^{\prime}}\mathrm{d}k^{\prime}\right\rangle\mathrm{d}k\mathrm{d}s$
@@
$T_{6,LH,L}^{x}=-2c_{\gamma}\int_{0}^{t}\iint_{|k-k^{\prime}|\leq\frac{|k|}{2},|k|<\nu^{-\frac{1}{3}}}\frac{\langle c\lambda_{k}s\rangle^{2J}}{M_{k}(s)}\langle k\rangle^{2m}\gamma_{k}\mathrm{Re}\langle\partial_{y}\psi_{k}\overline{\psi_{k-k^{\prime}}}+\psi_{k}\partial_{y}\overline{\psi_{k-k^{\prime}}},i(k-k^{\prime})\omega_{k^{\prime}}\rangle\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s.$
By Cauchy-Schwarz inequality in $y$,
$|T_{6,LH,L}^{x}|$ $\lesssim\int_{0}^{t}\iint_{\mathbb{R}^{2}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}(\|\partial_{y}\psi_{k}\|_{2}|k-k^{\prime}|\|\psi_{k-k^{\prime}}\|_{\infty}+|k|\|\psi_{k}\|_{2}\|\partial_{y}\psi_{k-k^{\prime}}\|_{\infty})\|\omega_{k^{\prime}}\|_{2}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\iint_{\mathbb{R}^{2}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}\langle c\lambda_{k^{\prime}}t\rangle^{J}\langle k^{\prime}\rangle^{m}\|\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s,$
Then using Young’s convolution inequality, Hölder’s inequality in $s$ and Lemma 3.5, we obtain
$|T_{6,LH,L}^{x}|$ $\lesssim\int_{0}^{t}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\nu^{-\frac{2}{3}}\gamma_{k}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\omega_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s$
$\lesssim\sup_{s\in[0,t]}(\nu^{-\frac{2}{3}}\mathcal{E}(s))^{\frac{1}{2}}\left(\int_{0}^{t}(\gamma_{k}\nu)^{-1}\widetilde{\mathcal{D}}\mathrm{d}s\right)^{\frac{1}{2}}\left(\int_{0}^{t}\left(\int_{\mathbb{R}}\|\nabla_{k}\psi_{k}\|_{\infty}\mathrm{d}k\right)^{2}\mathrm{d}s\right)^{\frac{1}{2}}$
$\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@
$|T_{6,LH,L}^{y}|$ $\lesssim\int_{0}^{t}\iint_{\mathbb{R}^{2}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\nu^{-\frac{1}{3}}\gamma_{k}^{\frac{1}{2}}\|\nabla_{k}\psi_{k}\|_{2}\langle c\lambda_{k^{\prime}}s\rangle^{J}\langle k^{\prime}\rangle^{m}\|\omega_{k^{\prime}}\|_{2}\|\nabla_{k-k^{\prime}}\psi_{k-k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\nu^{-\frac{7}{6}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$
@@
$|T_{6,HL,H}|$ $\lesssim\int_{0}^{t}\iint_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{J}\langle k\rangle^{m}\gamma_{k}\|\psi_{k}\|_{2}\langle c\lambda_{k-k^{\prime}}s\rangle^{J}\langle k-k^{\prime}\rangle^{m}\|\nabla_{k-k^{\prime}}^{\bot}\psi_{k-k^{\prime}}\|_{2}\|\nabla_{k^{\prime}}\omega_{k^{\prime}}\|_{\infty}\mathrm{d}k\mathrm{d}k^{\prime}\mathrm{d}s$
$\lesssim\int_{0}^{t}\left(\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\|\nabla_{k}\psi_{k}\|_{2}^{2}\mathrm{d}k\right)^{\frac{1}{2}}\int_{\mathbb{R}}\|\nabla_{k}\omega_{k}\|_{\infty}\mathrm{d}k\mathrm{d}s.$
@@
$\int_{|k|\geq\nu^{-\frac{1}{3}}}\langle c\lambda_{k}s\rangle^{2J}\langle k\rangle^{2m}\gamma_{k}^{2}\|\psi_{k}\|_{2}^{2}\mathrm{d}k\leq\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}}.$
Combining Lemma 3.4, we obtain
$|T_{6,HL,H}|\lesssim\int_{0}^{t}(\nu^{-\frac{1}{3}}\widetilde{\mathcal{D}})^{\frac{1}{2}}(\nu^{\frac{2}{3}}\mathcal{E})^{\frac{1}{2}}\nu^{-\frac{5}{6}}\widetilde{\mathcal{D}}^{\frac{1}{2}}\mathrm{d}s\lesssim\nu^{-\frac{2}{3}}\mathcal{D}(t)\sup_{s\in[0,t]}\mathcal{E}(s)^{\frac{1}{2}}.$